# The 2-systole on compact Kähler surfaces with positive scalar curvature

### v1

@@ Acknowledgment.
The author would first appreciate his advisor, Professor Gérard Besson, for introducing him to the field of PSC manifolds. The author also wants to thank Jian Wang for valuable comments on the manuscript.
@@ 2. Preliminaries
Let $(X^{n},\omega)$ be a compact Kähler manifold of complex dimension $n$. Locally we write $\omega=\sqrt{-1}\,g_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}$. We denote by $\mathcal{K}(X):=\{[\omega]\in H^{1,1}(X;\mathbb{R})\mid[\omega]>0\}$ the Kähler cone of $X$, which is an open convex cone in the finite-dimensional real vector space $H^{1,1}(X;\mathbb{R})$.
@@
$c_{1}(X)=\frac{1}{2\pi}[\operatorname{Ric}(\omega)]\in H^{1,1}(X)\cap H^{2}(X;\mathbb{R}).$
@@
depends only on the Kähler class $\omega]$. Clearly, $\min_{X}S(\omega)\leq\hat{S}([\omega])$, with equality if and only if $\omega$ is cscK.
@@
the $2$-systole. In contrast, the quantity that we will actually compute is the holomorphic $2$-systole introduced in Definition 1.3:
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]):=\inf\bigl\{[\omega]\cdot[C]\ \big|\ C\subset X\text{ an effective curve},\ 0\neq[C]\in H_{2}(X;\mathbb{Z})\bigr\}.$
Here we view an effective curve $C=\sum_{i}m_{i}C_{i}$ as an integral real 22–cycle and
$\omega]\cdot[C]:=\sum_{i}m_{i}\int_{C_{i}^{\mathrm{reg}}}\omega.$
In particular, every effective curve is a candidate for the $2$-systole, and we always have
@@
The basic scale–invariant functional that we will work with is
$\mathcal{J}_{X}([\omega]):=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\hat{S}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\frac{2n\pi\,c_{1}(X)\cup[\omega]^{n-1}}{[\omega]^{n}}.$
@@
for all $x\in K$ and all $\lambda>0$. Set $K_{1}:=\{x\in K\mid f(x)=1\}$. Assume that $K_{1}\neq\varnothing$. Then
@@
Let $(X,\omega)$ be a projective Kähler surface. Then
@@
    $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq\mathcal{J}_{X}([\omega])$, with equality if and only $\omega$ is cscK and $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic $1$-cycle;
@@
    $\sup_{\mathcal{K}^{+}(X)}|\mathcal{J}_{X}|<+\infty$, where $\mathcal{K}^{+}(X):=\left\{[\omega]\in\mathcal{K}(X);\penalty 10000\ c_{1}(X)\cdot[\omega]>0\right\}$
@@
We only prove (2) since (1) is obvious. Define $\mathcal{K}_{1}:=\left\{[\omega]\in\mathcal{K}^{+}(X);\penalty 10000\ \operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=1\right\}$, by Lemma 2.1, we have
$\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}=\sup_{\mathcal{K}_{1}}\mathcal{J}_{X}.$
We first show that $\overline{\mathcal{K}_{1}}\subset\mathcal{K}(X)$. Assume there is a sequence of Kähler metric $\omega_{\varepsilon}\rightarrow\omega_{0}\in\partial\mathcal{K}(X)$.
Claim: There exists an effective curve $C$, such that $\omega_{0}\cdot C=0$.
With the claim, we then have
$1=\operatorname{sys}^{\operatorname{hol}}_{2}(\omega_{\varepsilon})\leq[\omega_{\varepsilon}]\cdot[C]\rightarrow 0,$
which is a contradiction. Therefore, $\overline{\mathcal{K}_{1}}\subset\mathcal{K}(X)$.
Fix a Euclidean norm $\|\cdot\|$ on the finite dimensional vector space $V:=H^{1,1}(X)$, then $B=\{\alpha\in V;\penalty 10000\ \|\alpha\|=1\}$ is compact. Set $S_{1}=\overline{\mathcal{K}_{1}}\cap B$. Clearly, $S_{1}\subset\mathcal{K}(X)$ is also compact and the continuous function $\omega\mapsto\omega^{2}:=F(\omega)$ has minimum $m:=\min_{S_{1}}F$ and maximum $M:=\max_{S_{1}}F$ on $S_{1}$. For $\omega\in\mathcal{K}_{1}$, define $u=\omega/\|\omega\|\in S_{1}$, we then have $\omega^{2}=\|\omega\|^{2}\cdot u^{2}$. Thus,
$m\penalty 10000\ \|\omega\|^{2}\leq\omega^{2}\leq M\penalty 10000\ \|\omega\|^{2}.$
For any $\alpha,\beta\in V$, $\alpha\mapsto\beta\cdot\alpha$ is a continuous linear functional, so that
$|\beta\cdot\alpha|\leq C\penalty 10000\ \|\alpha\|.$
Consequently, we have
$\left|\frac{c_{1}(X)\cdot\omega}{\omega^{2}}\right|\leq\frac{C\penalty 10000\ \|\omega\|}{m\penalty 10000\ \|\omega\|^{2}}=\frac{C^{\prime}}{\|\omega\|}.$
It remains to prove $\|\omega\|$ is bounded away from 0 on $\mathcal{K}_{1}$. In particular, fix any effective curve $F$, $\omega\mapsto\omega\cdot F$ is a bounded linear functional on $\mathcal{K}_{1}$. So, we have
$1\leq\omega\cdot F\leq\tilde{C}\penalty 10000\ \|\omega\|,$
where $\tilde{C}>0$ only depends on $F$. This completes the proof. ∎
@@
Since $\omega_{0}$ is nef, we split into two cases.
Case 1: $\omega_{0}^{2}>0$: The claim follows from the Nakai-Moishezon criterion directly.
Case 2: $\omega_{0}^{2}=0$: Fix $t>3$ and set $D_{t}:=K_{X}+t\omega_{0}$. We claim that $D_{t}$ is nef.
By Mori’s cone Theorem (see, e.g. [Deb01, Theorem 7.49]) for smooth projective surfaces, There are countably many rational curves $C_{i}$ on $X$, satisfying $0<-K_{X}\cdot C_{i}\leq 3$, and
$\overline{\operatorname{NE}}(X)=\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}+\sum_{i}\mathbf{R}_{\geq 0}\left[C_{i}\right].$
For $\gamma\in\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}$ we have $D_{t}\cdot\gamma=K_{X}\cdot\gamma+t\,\omega_{0}\cdot\gamma\geq 0$ since $\omega_{0}$ is nef. For a $K_{X}$-negative extremal curve $C_{i}$, we have
$D_{t}\cdot C_{i}=K_{X}\cdot C_{i}+t\,\omega_{0}\cdot C_{i}\geq-3+t>0.$
@@
Since both $D_{t}$ and $\omega_{0}$ are nef, we have $D_{t}\cdot\omega_{0}\geq 0$. On the other hand, $\omega_{0}^{2}=0$ gives
$D_{t}\cdot\omega_{0}=(K_{X}+t\omega_{0})\cdot\omega_{0}=K_{X}\cdot\omega_{0}=-c_{1}(X)\cdot\omega_{0}<0,$
which is a contradiction. Therefore, Case 2 is ruled out. ∎
Blow-ups will play a central role in what follows. Let $\pi\colon\tilde{X}\to X$ be the blow-up of $X$ at a point $p$, with exceptional divisor $E$. Given a Kähler class $\omega]$ on $X$ we consider the family of classes on $\tilde{X}$
@@
$\varepsilon([\omega];p):=\sup\{\,t\geq 0\mid\pi^{*}[\omega]-t[E]\ \text{is nef}\,\}=\inf_{C\ni p}\frac{[\omega]\cdot[C]}{\mathrm{mult}_{p}(C)},$
where the infimum is taken over irreducible curves $C\subset X$ passing through $p$ and $\mathrm{mult}_{p}(C)$ denotes the multiplicity of $C$ at $p$. In particular, for every $0<t<\varepsilon([\omega];p)$ the class $\omega_{t}]$ is Kähler on $\tilde{X}$.
In this paper we will only consider Kähler classes $\omega]$ with $c_{1}(X)\cdot[\omega]>0$, since these are the classes that may contain PSC Kähler metrics. We denote the corresponding subcone by
Brown’s result [Bro24, Theorem A] shows that if $\omega]$ contains a PSC Kähler metric, then for all sufficiently small $t>0$ the classes $\omega_{t}]=\pi^{*}[\omega]-t[E]$ on the blow-up $\tilde{X}$ also contain PSC Kähler metrics. Therefore we may estimate $\mathcal{J}_{X}$ along such families and, ultimately, to obtain uniform bounds for $\mathcal{J}_{X}$ on PSC Kähler classes arising from $\mathbb{P}^{2}$ and ruled surfaces under finitely many blow-ups.

### v2

@@ Acknowledgment.
The author is deeply grateful to his advisor, Professor Gérard Besson, for introducing him to the study of positive scalar curvature and for his constant encouragement. The author also thanks Jian Wang for valuable comments on an earlier version of the manuscript.
@@ 2. Preliminaries
Let $(X^{n},\omega)$ be a compact Kähler manifold of complex dimension $n$. Locally we write
$\omega=\sqrt{-1}\,g_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}.$
We denote by
$H^{1,1}(X;\mathbb{R}):=H^{1,1}(X)\cap H^{2}(X;\mathbb{R})$
the space of real (1,1)(1,1)–classes, and by
$\mathcal{K}(X):=\{[\omega]\in H^{1,1}(X;\mathbb{R})\mid[\omega]>0\}$
the Kähler cone of $X$. It is an open convex cone in the finite-dimensional real vector space $H^{1,1}(X;\mathbb{R})$.
@@
$c_{1}(X)=\frac{1}{2\pi}[\operatorname{Ric}(\omega)]\in H^{1,1}(X;\mathbb{R}).$
@@
depends only on the Kähler class $\omega]$. Clearly,
$\min_{X}S(\omega)\leq\hat{S}([\omega]),$
with equality if and only if $\omega$ is cscK.
@@
the $2$-systole. In contrast, the quantity that we will actually compute is the following class-dependent holomorphic $2$-systole:
$\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]):=\inf\bigl\{[\omega]\cdot C\ \big|\ C\subset X\text{ an effective curve},\ [C]\neq 0\text{ in }H_{2}(X;\mathbb{Z})\bigr\}.$
Here an effective curve $C=\sum_{i}m_{i}C_{i}$ is viewed as an integral real 22–cycle, and we set
$\omega]\cdot C:=\sum_{i}m_{i}\int_{C_{i}^{\mathrm{reg}}}\omega.$
By abuse of notation, we do not distinguish a curve from its homology class when writing such pairings.
In particular, every effective curve determines a nontrivial integral 22–cycle, and therefore contributes a competitor in the definition of the $2$-systole. Hence we always have
@@
The basic scale-invariant functional that we will work with is the function
$\mathcal{J}_{X}:\mathcal{K}(X)\to\mathbb{R},\qquad\mathcal{J}_{X}([\omega]):=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\hat{S}([\omega]).$
Equivalently,
$\mathcal{J}_{X}([\omega])=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\frac{2n\pi\,c_{1}(X)\cup[\omega]^{n-1}}{[\omega]^{n}}.$
@@
In this paper we will only consider Kähler classes $\omega]$ with
$c_{1}(X)\cdot[\omega]>0,$
since these are the classes that may contain PSC Kähler metrics. We denote the corresponding subcone by
@@
for all $x\in K$ and all $\lambda>0$. Set
$K_{1}:=\{x\in K\mid f(x)=1\}.$
Assume that $K_{1}\neq\varnothing$. Then
@@
Let $(X,\omega)$ be a projective Kähler surface with $\mathcal{K}^{+}(X)\neq\emptyset$. Then:
@@
    For every $\omega]\in\mathcal{K}^{+}(X)$ and $\omega\in[\omega]$ with positive scalar curvature, we have $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq\mathcal{J}_{X}([\omega])$, with equality if and only if $\omega$ is cscK and $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic $1$-cycle;
@@
    $\sup_{[\omega]\in\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])<+\infty$.
@@
Part (1) follows immediately from
$\min_{X}S(\omega)\leq\hat{S}([\omega])\quad\text{and}\quad\operatorname{sys}_{2}(\omega)\leq\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]),$
with equality precisely when $\omega$ is cscK and $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic $1$-cycle. We therefore prove only Part (2).
Set $\mathcal{K}_{1}:=\{[\omega]\in\mathcal{K}^{+}(X)\mid\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=1\}$. By Lemma 2.1,
$\sup_{[\omega]\in\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])=\sup_{[\omega]\in\mathcal{K}_{1}}\mathcal{J}_{X}([\omega]).$
We first show that $\overline{\mathcal{K}_{1}}\subset\mathcal{K}(X)$. Assume for contradiction that there exists a sequence of Kähler classes $\omega_{\varepsilon}]\in\mathcal{K}_{1}$ such that $\omega_{\varepsilon}]\to[\omega_{0}]\in\partial\mathcal{K}(X)$.
Claim. There exists an effective curve $C\subset X$ such that
$\omega_{0}]\cdot C=0.$
Granting the claim, we obtain
$1=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{\varepsilon}])\leq[\omega_{\varepsilon}]\cdot C\to[\omega_{0}]\cdot C=0,$
a contradiction. Hence $\overline{\mathcal{K}_{1}}\subset\mathcal{K}(X)$.
Now fix a Euclidean norm $\|\cdot\|$ on the finite-dimensional real vector space $V:=H^{1,1}(X;\mathbb{R})$, and let $B:=\{\alpha\in V\mid\|\alpha\|=1\}$. Set $S_{1}:=\overline{\mathcal{K}_{1}}\cap B$. Then $S_{1}\subset\mathcal{K}(X)$ is compact. Consider the continuous function $F:\alpha\mapsto\alpha^{2}$ on $V$. Since $S_{1}\subset\mathcal{K}(X)$, we have $F>0$ on $S_{1}$, and therefore
$m:=\min_{S_{1}}F>0,\qquad M:=\max_{S_{1}}F<\infty.$
For any $\alpha\in\mathcal{K}_{1}$, define
$u:=\frac{\alpha}{\|\alpha\|}\in S_{1}.$
Then $\alpha^{2}=\|\alpha\|^{2}u^{2}$, so
$m\,\|\alpha\|^{2}\leq\alpha^{2}\leq M\,\|\alpha\|^{2}.$
Since $\alpha\mapsto c_{1}(X)\cdot\alpha$ is a continuous linear functional on $V$, there exists $C>0$ such that
$|c_{1}(X)\cdot\alpha|\leq C\|\alpha|\qquad\text{for all }\alpha\in V.$
Hence
$\left|\frac{c_{1}(X)\cdot\alpha}{\alpha^{2}}\right|\leq\frac{C\|\alpha\|}{m\|\alpha\|^{2}}=\frac{C^{\prime}}{\|\alpha\|}.$
It remains to show that $\|\alpha\|$ is bounded away from 0 on $\mathcal{K}_{1}$. Take any effective curve $F$. Since $\alpha\mapsto\alpha\cdot F$ is a continuous linear functional on $V$, there exists $\widetilde{C}>0$ such that
$|\alpha\cdot F|\leq\widetilde{C}\,\|\alpha\|\qquad\text{for all }\alpha\in V.$
On the other hand, for every $\alpha\in\mathcal{K}_{1}$, we have
$1=\operatorname{sys}^{\operatorname{hol}}_{2}(\alpha)\leq\alpha\cdot F.$
Therefore
$1\leq\alpha\cdot F\leq\widetilde{C}\,\|\alpha\|,$
so $\|\alpha\|\geq\widetilde{C}^{-1}$ on $\mathcal{K}_{1}$. This proves Part (2). ∎
@@
Set $\alpha_{0}:=[\omega_{0}]\in\partial\mathcal{K}(X)$. Since $\alpha_{0}$ is nef, we split into two cases.
Case 1: $\alpha_{0}^{2}>0$. If $\alpha_{0}\cdot C>0$ for every irreducible curve $C\subset X$, then $\alpha_{0}$ is ample by the Nakai–Moishezon criterion for surfaces, contradicting $\alpha_{0}\in\partial\mathcal{K}(X)$. Hence there exists an irreducible curve $C\subset X$ such that
$\alpha_{0}\cdot C=0.$
Case 2: $\alpha_{0}^{2}=0$. We first show that $c_{1}(X)\cdot\alpha_{0}=0$. Assume, for contradiction, that $c_{1}(X)\cdot\alpha_{0}>0$. Fix $t>3$ and set
$D_{t}:=K_{X}+t\alpha_{0}.$
We claim that $D_{t}$ is nef.
By the Mori’s Cone Theorem for smooth projective surfaces (see, for instance, [9, Theorem 7.49]), there exist countably many rational curves $C_{i}$ on $X$ such that
$0<-K_{X}\cdot C_{i}\leq 3,$
and
$\overline{\operatorname{NE}}(X)=\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}+\sum_{i}\mathbb{R}_{\geq 0}[C_{i}],$
where
$\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}:=\{\gamma\in\overline{\operatorname{NE}}(X)\mid K_{X}\cdot\gamma\geq 0\}.$
For $\gamma\in\overline{\operatorname{NE}}(X)_{K_{X}\geq 0}$ we have
$D_{t}\cdot\gamma=K_{X}\cdot\gamma+t\,\alpha_{0}\cdot\gamma\geq 0$
since $\alpha_{0}$ is nef. For a $K_{X}$-negative extremal curve $C_{i}$, we have
$D_{t}\cdot C_{i}=K_{X}\cdot C_{i}+t\,\alpha_{0}\cdot C_{i}\geq-3+t>0.$
@@
Since both $D_{t}$ and $\alpha_{0}$ are nef, we have $D_{t}\cdot\alpha_{0}\geq 0$. On the other hand, $\alpha_{0}^{2}=0$ gives
$0\leq D_{t}\cdot\alpha_{0}=(K_{X}+t\alpha_{0})\cdot\alpha_{0}=K_{X}\cdot\alpha_{0}=-c_{1}(X)\cdot\alpha_{0}<0,$
which is a contradiction. Therefore, $c_{1}(X)\cdot\alpha_{0}=0$.
Since $\alpha_{0}\neq 0$ is nef with $\alpha_{0}^{2}=0$, the Hodge index theorem implies that the intersection form is non-positive on $\alpha_{0}^{\perp}$, and
$D_{t}\cdot\alpha_{0}=(K_{X}+t\alpha_{0})\cdot\alpha_{0}=K_{X}\cdot\alpha_{0}=-c_{1}(X)\cdot\alpha_{0}=0,$
we obtain $D_{t}^{2}\leq 0$. On the other hand, $D_{t}$ is nef, hence $D_{t}^{2}\geq 0$. Therefore $D_{t}^{2}=0$. The equality case in the Hodge index theorem now implies that $D_{t}$ is numerically proportional to $\alpha_{0}$, namely $D_{t}\equiv\lambda_{t}\,\alpha_{0}$ for some $\lambda_{t}\in\mathbb{R}$. Hence
$K_{X}\equiv a\,\alpha_{0}$
for some $a\in\mathbb{R}$. Since $\mathcal{K}^{+}(X)\neq\emptyset$, we may choose $\beta\in\mathcal{K}^{+}(X)$. Then
$(-K_{X})\cdot\beta=c_{1}(X)\cdot\beta>0,$
while $\alpha_{0}\cdot\beta>0$ because $\alpha_{0}$ is nef and nonzero. Therefore $a>0$. In particular, $X$ admits a Kähler class of positive total scalar curvature. By [15], $X$ is uniruled. Hence $\kappa(X)=$-$\infty$. By the Enriques–Kodaira classification of projective surfaces, the minimal model of $X$ is therefore either rational or ruled (see, for example, [2, Chapter V]).
If the minimal model of $X$ is ruled over a curve of genus $g$, then by the canonical divisor formula for ruled surfaces
$K_{X}^{2}=8(1-g)$
on the minimal model (see [14, Chapter V, §2]). Since each blow-up decreases $K^{2}$ by $1$, and since here
$K_{X}^{2}=(-K_{X})^{2}=a^{2}\alpha_{0}^{2}=0,$
it follows that necessarily $g=1$ and $X$ is already minimal. Thus the ruled case reduces to a minimal ruled surface over an elliptic curve. We now treat the two possibilities separately.
Subcase 2a: $X$ is rational. Since $X$ is rational, $\chi(\mathcal{O}_{X})=1$. Applying Riemann–Roch theorem to $D=-K_{X}$, we get
$\chi(\mathcal{O}_{X}(-K_{X}))=\chi(\mathcal{O}_{X})+\frac{(-K_{X})^{2}-K_{X}\cdot(-K_{X})}{2}=1+\frac{2K_{X}^{2}}{2}=1.$
By Serre duality,
$h^{2}(X,-K_{X})=h^{0}(X,2K_{X})=0,$
since $\kappa(X)=$-$\infty$. Hence
$h^{0}(X,-K_{X})-h^{1}(X,-K_{X})=1,$
and in particular
$h^{0}(X,-K_{X})\geq 1.$
Thus there exists a nonzero effective divisor $D\in|-K_{X}|$. Moreover,
$\alpha_{0}\cdot D=\frac{1}{a}(-K_{X})\cdot D=\frac{1}{a}(-K_{X})^{2}=0.$
Write $D=\sum_{j}m_{j}D_{j}$ as a sum of irreducible components. Since $\alpha_{0}$ is nef, each $\alpha_{0}\cdot D_{j}\geq 0$, and the equality $\alpha_{0}\cdot D=0$ forces $\alpha_{0}\cdot D_{j}=0$ for every component $D_{j}$. Thus any irreducible component of $D$ gives the desired curve.
Subcase 2b: $X$ is ruled over an elliptic curve. Since $X$ is a minimal ruled surface over an elliptic curve, let $C_{0}$ be a section of minimal self-intersection and $F$ be a fibre (as in Section 4). Then
$C_{0}^{2}=-e,\qquad e\geq-1,\qquad-K_{X}\equiv 2C_{0}+eF.$
Because $K_{X}$ is nef, we have
$0\leq(-K_{X})\cdot C_{0}=(2C_{0}+eF)\cdot C_{0}=2(-e)+e=-e.$
Hence $e=0$ or 1-1.
If $e=0$, then $K_{X}\equiv 2C_{0}\equiv a\,\alpha_{0}$, and therefore
$\alpha_{0}\cdot C_{0}=\frac{1}{a}(-K_{X})\cdot C_{0}=\frac{2}{a}\,C_{0}^{2}=0.$
Thus $C_{0}$ is the required irreducible effective curve.
If e=1e=-1, then $K_{X}\equiv 2C_{0}-F$. By [11, Proposition 3.2] (see also [16, Corollary 2.2]), there exists a smooth elliptic curve $E\subset X$ such that $E\equiv 2C_{0}-F$. Hence
$\alpha_{0}\cdot E=\frac{1}{a}(-K_{X})\cdot E=\frac{1}{a}(-K_{X})^{2}=0.$
Thus $E$ is the desired irreducible effective curve. This completes the proof in Case 2. ∎
Blow-ups will play a central role in what follows. Let $\pi\colon\tilde{X}\to X$ be the blow-up of $X$ at a point $p$, with exceptional divisor $E$. Given a Kähler class $\omega]$ on $X$, we consider the family of classes on $\tilde{X}$
@@
$\varepsilon([\omega];p):=\sup\{\,t\geq 0\mid\pi^{*}[\omega]-t[E]\ \text{is nef}\,\}=\inf_{C\ni p}\frac{[\omega]\cdot C}{\mathrm{mult}_{p}(C)},$
where the infimum is taken over irreducible curves $C\subset X$ passing through $p$, and $\mathrm{mult}_{p}(C)$ denotes the multiplicity of $C$ at $p$. In particular, for every $0<t<\varepsilon([\omega];p)$, the class $\omega_{t}]$ is Kähler on $\tilde{X}$.
Brown’s result [8, Theorem A] shows that if $\omega]$ contains a PSC Kähler metric, then for all sufficiently small $t>0$ the classes
$\omega_{t}]=\pi^{*}[\omega]-t[E]$
on the blow-up $\tilde{X}$ also contain PSC Kähler metrics. Therefore we may estimate $\mathcal{J}_{X}$ along such families and, ultimately, obtain uniform bounds for $\mathcal{J}_{X}$ on PSC Kähler classes arising from $\mathbb{P}^{2}$ and ruled surfaces under finitely many blow-ups.